OpenStudy (anonymous):

Can anyone help me with this topic? I know this is pretty easy but, how do you factor trinomials?! :|

OpenStudy (anonymous):

depends on the specific equation.

OpenStudy (anonymous):

i mean binomials, a^3x + 125

OpenStudy (anonymous):

did you type that correctly? use the equation tool if you need to.

OpenStudy (kira_yamato):

\[(x-y)(x+y) = x^2 - y^2\]

OpenStudy (anonymous):

\[a ^3^x + 125\]

OpenStudy (anonymous):

Okay fail, the exponent is 3x

OpenStudy (kira_yamato):

\[a^{3x}-125?\]

OpenStudy (anonymous):

Yes :D Except that the operation is a plus :)

OpenStudy (anonymous):

so a^3x-125= (a^(3x/2)-5sqrt5)(a^(3x/2)+5sqrt5)

OpenStudy (anonymous):

if it is +125, then it will have an i in it

OpenStudy (anonymous):

5sqrt5i

OpenStudy (anonymous):

Oohh.. Okay okay, im getting it :)

OpenStudy (anonymous):

Thanks! :D